# WMCopier: Forging Invisible Image Watermarks on Arbitrary Images

### v2

@@ Abstract
Invisible Image Watermarking is crucial for ensuring content provenance and accountability in generative AI. While Gen-AI providers are increasingly integrating invisible watermarking systems, the robustness of these schemes against forgery attacks remains poorly characterized. This is critical, as forging traceable watermarks onto illicit content leads to false attribution, potentially harming the reputation and legal standing of Gen-AI service providers who are not responsible for the content. In this work, we propose WMCopier, an effective watermark forgery attack that operates without requiring any prior knowledge of or access to the target watermarking algorithm. Our approach first models the target watermark distribution using an unconditional diffusion model, and then seamlessly embeds the target watermark into a non-watermarked image via a shallow inversion process. We also incorporate an iterative optimization procedure that refines the reconstructed image to further trade off the fidelity and forgery efficiency. Experimental results demonstrate that WMCopier effectively deceives both open-source and closed-source watermark systems (e.g., Amazon’s system), achieving a significantly higher success rate than existing methods [^1] We have reported this to Amazon AGI’s Responsible AI team and collaborated on developing potential defense strategies. For the official statement from Amazon, see the “Broader Impact” in Appendix I.. Additionally, we evaluate the robustness of forged samples and discuss the potential defenses against our attack. Code is available at: https://anonymous.4open.science/r/WMCopier-E752.
@@ 1 Introduction
Existing watermark forgery attacks are broadly categorized into two scenarios: the black-box setting and the no-box setting. In the black-box setting, the attacker has partial access to the watermarking system: such as knowledge of the specific watermarking algorithm [31], the ability to obtain paired data (clean images and their watermark versions) via the embedding interface [35, 44], and query access to the watermark detector [31]. However, such black-box access is unrealistic in practice, as the watermark embedding process is typically integrated into the generative service itself, rendering it inaccessible to end users, thus disabling paired data acquisition. Moreover, service providers rarely disclose the specific watermarking algorithms they employ [10]. Therefore, our focus is primarily on the no-box setting, where the attacker lacks both knowledge of the watermarking algorithm and access to its implementation. The available information is only a collection of generated images with unknown watermarking algorithms. Under this setting, Yang et al.[47] attempt to extract the watermark pattern by computing the mean residual between watermarked images and natural images from ImageNet[34]. The extracted watermark pattern is then directly added to forged images at the pixel level. However, this coarse approximation limits the effectiveness of the forgery, as the natural images from ImageNet inevitably differ from the clean counterparts of the watermarked images.
@@
To evaluate the effectiveness of WMCopier, we perform comprehensive experiments across a range of watermarking schemes, including a closed-source one (Amazon’s system). Experimental results demonstrate that our attack achieves a high forgery success rate while preserving excellent visual fidelity. Furthermore, we conduct a comparative robustness analysis between genuine and forged watermarks. Finally, we explore potential defense strategies, including a multi-message strategy and semantic watermarking, which provide practical guidance for improving future watermark design and deployment.
@@
  • We explore potential defense strategies, including a multi-message strategy and semantic watermarking, that provide insights to improve future watermarking systems.
@@ 2.2 Invisible Image Watermarking
$V(m,m^{\prime},\rho)=\begin{cases}\text{Watermarked},&\text{if }\texttt{Bit-Accuracy}(m,m^{\prime})\geq\rho;\\ \text{Non-Watermarked},&\text{otherwise.}\end{cases}$ (5)
@@ 3 Threat Model
    The attacker does not know the target watermarking scheme or its internal parameters. They have no access to embed watermarks into their own images or the corresponding detection pipeline.
@@
    The attacker can collect a subset of watermarked images from AI-generated content platforms (e.g., PromptBase [41], PromptHero [42]) or directly query the target Gen-AI service.
@@ 4.1 Watermark Estimation
Now consider a clean image $x$ and its watermarked version $x^{w}=x+w$, where $w$ denotes the embedded watermark signal, which can also be interpreted as the perturbation introduced by the embedding process. During the forward diffusion process, we have:
@@
Figure 2: Watermark detectability of four open-source watermarking schemes throughout the diffusion and denoising processes ($T=1000$). As a reference, the bit accuracy of non-watermarked images remains around 50%.
@@ 4.2 Watermark Injection
Therefore, we propose a shallow inversion strategy that performs the inversion process up to an early timestep $T_{S}\ll T$. By skipping deeper diffusion steps that contribute minimally to watermark injection yet substantially distort image semantics, our method effectively preserves the visual fidelity of reconstructed images while ensuring reliable watermark injection.
@@ 5 Evaluation
Attack Parameters and Baselines. For the diffusion model, we adopt DDIM sampling DDIM sampling with a total step $T=100$ and perform inversion up to step $T_{S}=40$. Further details regarding the training of the diffusion model are provided in the Appendix G. For the refinement procedure, we set the trade-off coefficient $\lambda$ as 100, the number of refinement iterations $L$ as 100, a low-noise step $t_{l}$ in the refinement as 1 and the step size $\eta$ as $1\times 10^{-4}$ by default. To balance the attack performance and the potential cost of acquiring generated images (e.g., fees from GenAI services), we set the size of the auxiliary dataset $\mathcal{D}_{\text{aux}}$ to 5,000 in our main experiments. For comparison, we consider the method by Yang et al. [47] that operates under the same no-box setting as ours, and Wang et al. [44] that assumes a black-box setting with access to paired watermarked and clean images.
Metrics. We evaluate the visual quality of forged images using Peak Signal-to-Noise Ratio (PSNR), defined as $\mathrm{PSNR}(x,\hat{x^{f}})=-10\cdot\log_{10}\left(\mathrm{MSE}(x,\hat{x^{f}})\right),$ where $x$ is the clean image and $\hat{x_{f}}$ is the forged image after the refinement process. A higher PSNR indicates better visual fidelity, i.e., the forged image is more similar to the original. We evaluate the attack effectiveness in terms of bit accuracy and false positive rate (FPR). Bit accuracy measures the proportion of watermark bits in the extracted message that match the target. FPR refers to the rate at which forged samples are incorrectly identified as valid watermarked images. A higher FPR thus indicates a more successful attack. Following common practice [28], we report FPR at a threshold calibrated to yield a 5% false positive rate on clean images.
@@
Watermark scheme Dataset PSNR↑ Forged Bit-acc↑ FPR@0.05↑ PSNR↑ Forged Bit-acc.↑ FPR@0.05↑ PSNR↑ Forged Bit-acc.↑ FPR@0.05↑
DWT-DCT MS-COCO 31.33 74.32% 81.10% 32.87 53.08% 8.00% 33.69 89.19% 90.80%
CelebAHQ 32.19 81.29% 85.60% 32.90 53.68% 11.00% 35.29 89.46% 95.20%
ImageNet 30.16 79.64% 83.30% 32.92 51.96% 6.40% 33.75 88.25% 89.20%
Diffusiondb 31.87 78.22% 82.80% 32.90 51.59% 6.60% 33.84 85.17% 86.00%
HiddeN MS-COCO 31.02 80.56% 99.30% 29.68 63.12% 32.80% 31.74 99.34% 100.00%
CelebAHQ 31.57 82.28% 99.90% 29.79 61.52% 21.50% 33.12 98.08% 99.90%
ImageNet 31.24 78.61% 95.00% 29.78 62.66% 32.90% 31.76 98.99% 99.90%
Diffusiondb 30.74 79.99% 99.20% 29.68 63.36% 35.40% 31.46 98.83% 99.80%
RivaGAN MS-COCO 32.94 93.26% 100.00% 29.12 50.80% 3.00% 34.07 95.74% 100.00%
CelebAHQ 32.64 93.67% 100.00% 29.23 52.29% 2.90% 35.28 98.61% 100.00%
ImageNet 33.11 90.94% 99.60% 29.22 50.92% 2.70% 33.87 93.83% 100.00%
Diffusiondb 33.31 89.69% 96.70% 29.12 48.70% 1.10% 34.50 90.43% 97.40%
Stable Signature MS-COCO 28.87 91.68% 98.40% 30.77 52.67% 5.90% 31.29 98.04% 99.90%
CelebAHQ 32.33 79.90% 92.40% 30.51 51.73% 1.50% 30.54 96.04% 100.00%
ImageNet 29.59 85.77% 89.90% 30.75 51.59% 3.80% 31.33 97.03% 100.00%
Diffusiondb 31.11 89.24% 98.10% 30.65 52.69% 4.60% 31.59 96.24% 99.80%
Average 31.50 84.32% 93.83% 30.62 54.52% 11.26% 32.94 94.58% 97.37%
@@ 5.2 Attacks on Closed-Source Watermarking Systems
In this subsection, we evaluate the effectiveness of our attack and Yang’s method in attacking the Amazon watermarking scheme. The results are shown in Table 2. The success rate (SR), which represents the proportion of images detected as watermarked, and the confidence levels (Con.) returned by the API, are used to evaluate the effectiveness of the attacks on deployed watermarking systems.
Compared with Yang’s method, our attack achieves superior performance in terms of both visual fidelity and forgery effectiveness. Specifically, our method achieves an average PSNR exceeding 30dB and a success rate(SR) close to 100%, whereas Yang’s method typically results in PSNR values below 25dB and SR ranging from 28% to 42%.
Furthermore, our forged images generally receive a confidence level of 3—the highest rating defined by Amazon’s watermark detection API—while Yang’s results consistently remain at level 2. Since Amazon does not disclose the exact computation of the confidence score, we guess that it may correlate with bit accuracy, based on common assumptions [28]. To further investigate this, we analyzed the distribution of forged bit accuracy of both our method and Yang’s on open-source watermarking schemes. As shown in Figure 4, our method achieves 80%–100% bit accuracy on RivaGan, significantly outperforming Yang’s method, which remains below 70%.
@@ 5.3 Ablation Study
To evaluate the impact of parameter choices on image quality and forgery effectiveness, we conduct two sets of ablation studies by varying (i) the number of refinement optimization steps $L$ and (ii) the trade-off coefficient $\lambda$. As shown in Figure 5, increasing $L$ initially improves both PSNR and forged bit accuracy, with performance saturating beyond $L=100$. In contrast, larger $\lambda$ values continuously enhance PSNR but lead to a slight degradation in bit accuracy, likely due to over-regularization. Given that PSNR values above 30 dB are generally considered visually indistinguishable to the human eye, we adopt $\lambda=100$ in our main experiments. The results presented in Table 9 in Appendix F, further validate the effectiveness of the refinement.
@@ 6.1 Image Watermarking
Post-processing methods embed watermark messages into images after generation. Classical approaches (e.g., LSB [7], DWT-DCT [1, 32]) suffer from poor robustness under common distortions such as compression and noise. Neural approaches mitigate these issues by combining encoder-decoder architectures and adversarial training [55, 40, 13, 20]. However, these methods often rely on heavy training and may generalize poorly to unknown attacks.
In-processing methods embed watermarks during image generation, either by modifying training data or model weights [50, 49, 28], or by adjusting specific components such as diffusion decoders [14]. Recent trends explore semantic watermarking, which binds messages to generative semantics (e.g., Tree-Ring [46]; Gaussian shading [48]). However, semantic watermarking remains in an early stage and has not seen real-world deployment [31], and thus falls beyond the scope of this work.
@@ 6.2 Watermark Forgery
Prior studies on watermark forgery primarily operate under white-box [24] or black-box [35, 44, 25, 31] assumptions, where the attacker either has access to the watermarking model or can embed the watermark in their images. These approaches rely on strong assumptions that may not hold in real-world scenarios. In contrast, the no-box setting assumes that only watermarked images are accessible. Yang et al. [47] propose a heuristic approach in this setting, wherein they estimate the watermark signal by averaging the residuals between watermarked and clean images, and subsequently embed the estimated pattern directly at the pixel level. This is the setting we focus on in this work, as it more realistically reflects practical constraints.
@@ 7 Defense Analysis
In this section, we analyze potential defenses against our WMCopier, including a multi-message strategy and semantic watermarking. We provide detailed experimental evaluations of these defenses in the Appendix E.

### v3

@@ Abstract
Invisible Image Watermarking is crucial for ensuring content provenance and accountability in generative AI. While Gen-AI providers are increasingly integrating invisible watermarking systems, the robustness of these schemes against forgery attacks remains poorly characterized. This is critical, as forging traceable watermarks onto illicit content leads to false attribution, potentially harming the reputation and legal standing of Gen-AI service providers who are not responsible for the content. In this work, we propose WMCopier, an effective watermark forgery attack that operates without requiring any prior knowledge of or access to the target watermarking algorithm. Our approach first models the target watermark distribution using an unconditional diffusion model, and then seamlessly embeds the target watermark into a non-watermarked image via a shallow inversion process. We also incorporate an iterative optimization procedure that refines the reconstructed image to further trade off the fidelity and forgery efficiency. Experimental results demonstrate that WMCopier effectively deceives both open-source and closed-source watermark systems (e.g., Amazon’s system), achieving a significantly higher success rate than existing methods [^1] We have reported this to Amazon AGI’s Responsible AI team and collaborated on developing potential defense strategies. For the official statement from Amazon, see Appendix H.. Additionally, we evaluate the robustness of forged samples and discuss the potential defenses against our attack. Code is available at: https://github.com/holdrain/WMCopier.
@@ 1 Introduction
Existing watermark forgery attacks are broadly categorized into two scenarios: the black-box setting and the no-box setting. In the black-box setting, the attacker has partial access to the watermarking system: such as knowledge of the specific watermarking algorithm [14], the ability to obtain paired data (clean images and their watermark versions) via the embedding interface [15, 16], or query access to the watermark detector [14]. However, such black-box access is unrealistic in practice, as the watermark embedding process is typically integrated into the generative service itself, rendering it inaccessible to end users, thus disabling paired data acquisition. Moreover, service providers rarely disclose the specific watermarking algorithms they employ [5]. Therefore, our focus is primarily on the no-box setting, where the attacker has neither knowledge of the watermarking algorithm nor access to its implementation, and only a collection of generated images with unknown watermarking schemes is available. Under this setting, Yang et al. [10] attempt to extract the watermark pattern by computing the mean residual between watermarked images and natural images from ImageNet [17], and then directly adding the estimated pattern to forged images at the pixel level. However, this achieves limited performance because it assumes that the watermark signal remains constant across all images. Moreover, its estimation is further hindered by the domain gap between ImageNet images and the unknown clean counterparts of the watermarked samples.
@@
To evaluate the effectiveness of WMCopier, we perform comprehensive experiments across a range of watermarking schemes, including a closed-source one (Amazon’s system). Experimental results demonstrate that our attack achieves a high forgery success rate while preserving excellent visual fidelity. Furthermore, we conduct a comparative robustness analysis between genuine and forged watermarks. Finally, we explore a multi-message defense strategy that provides practical guidance for improving future watermark design and deployment.
@@
  • We explore a potential defense strategy that provides insights to improve future watermarking systems.
@@ 2.2 Invisible Image Watermarking
[…]
@@ 3 Threat Model
    The attacker does not know the target watermarking scheme and its internal parameters. They have no access to embed watermarks into their own images and the corresponding detection pipeline.
@@
    The attacker can collect a subset of watermarked images from AI-generated content platforms (e.g., PromptBase [29], PromptHero [30]) or directly use the target Gen-AI service.
@@ 4.1 Watermark Estimation
Now consider a clean image $x$ and its watermarked version $x^{w}=x+w(w)$, where $w$ denotes the embedded watermark signal, which can also be interpreted as the perturbation introduced by the embedding process. During the forward diffusion process, we have:
@@
Figure 2: Watermark detectability of four open-source watermarking schemes throughout the diffusion and denoising processes ($T=1000$). As a reference, the bit accuracy of non-watermarked images remains around 0.5.
@@ 4.2 Watermark Injection
Therefore, we propose a shallow inversion strategy that performs the inversion process up to an early timestep $T_{S}<T$. By skipping deeper diffusion steps that contribute minimally to watermark injection yet substantially distort image semantics, our method effectively preserves the visual fidelity of reconstructed images while ensuring reliable watermark injection.
@@ 5 Evaluation
Attack Parameters and Baselines. For the diffusion model, we adopt DDIM sampling DDIM sampling with a total step $T=100$ and perform inversion up to step $T_{S}=40$. Further details regarding the training of the diffusion model are provided in the Appendix F. For the refinement procedure, we set the trade-off coefficient $\lambda$ as 100, the number of refinement iterations $L$ as 100, a low-noise step $t_{l}$ in the refinement as 1 and the step size $\eta$ as $1\times 10^{-4}$ by default. To balance the attack performance and the potential cost of acquiring generated images (e.g., fees from GenAI services), we set the size of the auxiliary dataset $\mathcal{D}_{\text{aux}}$ to 5,000 in our main experiments. For comparison, we consider the method by Yang et al. [10] that operates under the same no-box setting as ours, and Wang et al. [16] that assumes a black-box setting with access to paired watermarked and clean images.
[…]
@@
[…]
DWT-DCT MS-COCO 31.33 74.32% 57.20% 32.87 53.08% 0.50% 33.69 89.19% 60.20%
CelebAHQ 32.19 81.29% 50.70% 32.90 53.68% 0.10% 35.29 89.46% 53.20%
ImageNet 30.16 79.64% 55.10% 32.92 51.96% 0.20% 33.75 88.25% 55.80%
Diffusiondb 31.87 78.22% 50.80% 32.90 51.59% 0.40% 33.84 85.17% 54.30%
HiddeN MS-COCO 31.02 80.56% 77.60% 29.68 63.12% 0.00% 31.74 99.34% 95.90%
CelebAHQ 31.57 82.28% 80.20% 29.79 61.52% 0.00% 33.12 98.08% 92.50%
ImageNet 31.24 78.61% 83.90% 29.78 62.66% 0.00% 31.76 98.99% 94.30%
Diffusiondb 30.74 79.99% 79.20% 29.68 63.36% 0.00% 31.46 98.83% 94.60%
RivaGAN MS-COCO 32.94 93.26% 88.80% 29.12 50.80% 0.00% 34.07 95.74% 90.90%
CelebAHQ 32.64 93.67% 93.80% 29.23 52.29% 0.00% 35.28 98.61% 96.00%
ImageNet 33.11 90.94% 71.40% 29.22 50.92% 0.00% 33.87 93.83% 77.10%
Diffusiondb 33.31 89.69% 80.60% 29.12 48.70% 0.00% 34.50 90.43% 84.80%
Stable Signature MS-COCO 28.87 91.68% 88.90% 30.77 52.67% 0.00% 31.29 98.04% 94.60%
CelebAHQ 32.33 79.90% 90.10% 30.51 51.73% 0.00% 30.54 96.04% 100.00%
ImageNet 29.59 85.77% 85.90% 30.75 51.59% 0.00% 31.33 97.03% 98.60%
Diffusiondb 31.11 89.24% 92.10% 30.65 52.69% 0.00% 31.59 96.24% 96.60%
Average 31.50 84.32% 76.64% 30.62 54.52% 0.08% 32.94 94.58% 83.71%
@@ 5.2 Attacks on Closed-Source Watermarking Systems
In this subsection, we evaluate the effectiveness of our attack and Yang’s method in attacking the Amazon watermarking scheme. The results are shown in Table 2. The success rate (SR), which represents the proportion of images detected as watermarked, and the confidence levels (Con.) returned by the API, are used to evaluate the effectiveness of the attacks on deployed watermarking systems. Compared with Yang’s method, our attack achieves superior performance in terms of both visual fidelity and forgery effectiveness. Specifically, our method achieves an average PSNR exceeding 30dB and a success rate(SR) close to 100%, whereas Yang’s method typically results in PSNR values below 25dB and SR ranging from 28% to 42%.
Furthermore, our forged images generally receive a confidence level of 3—the highest rating defined by Amazon’s watermark detection API—while Yang’s results consistently remain at level 2. Since Amazon does not disclose the exact computation of the confidence score, we guess that it may correlate with bit accuracy, based on common assumptions [28]. To further investigate this, we analyzed the distribution of forged bit accuracy of both our method and Yang’s on a open-source watermarking scheme. As shown in Figure 4, our method achieves over 80% bit accuracy on RivaGan, significantly outperforming Yang’s method, which remains below 70%.
@@ 5.3 Ablation Study
To evaluate the impact of parameter choices on image quality and forgery effectiveness, we conduct two sets of ablation studies by varying (i) the number of refinement optimization steps $L$ and (ii) the trade-off coefficient $\lambda$. As shown in Figure 5, increasing $L$ initially improves both PSNR and forged bit accuracy, with performance saturating beyond $L=100$. In contrast, larger $\lambda$ values continuously enhance PSNR but lead to a slight degradation in bit accuracy, likely due to over-regularization. While higher PSNR values generally indicate better visual fidelity, we note that visible artefacts may still occur even at elevated PSNR levels. Nevertheless, since an attacker may prioritize forgery success over perceptual quality, we adopt $\lambda=100$ in our main experiments. The results presented in Table 11 in Appendix E further validate the effectiveness of the refinement process.
@@ 6.1 Image Watermarking
Post-processing methods embed watermark messages into images after generation. Non-learning-based methods (e.g., LSB [43], DWT-DCT [40, 44]) suffer from poor robustness under common distortions such as compression and noise. Neural network-based approaches mitigate these issues by combining encoder-decoder architectures and adversarial training [41, 45, 46, 47]. However, these methods often rely on heavy training and may generalize poorly to unknown attacks.
In-processing methods embed watermarks during image generation, either by modifying training data or model weights [19, 48, 28], or by adjusting specific components such as diffusion decoders [26]. Recent trends explore semantic watermarking, which binds messages to generative semantics (e.g., Tree-Ring [49]; Gaussian shading [50]). However, semantic watermarking has not seen real-world deployment [14]. We discuss the effectiveness of our attack on the semantic watermarking in the Appendix D.2.
@@ 6.2 Watermark Forgery
Kutter et al. [51] first introduced the concept, also known as the watermark copy attack, under the assumption that the watermark signal was a fixed constant. While this assumption was reasonable for early handcrafted watermarking methods, it no longer holds for modern neural network-based schemes. Subsequent studies [52, 16, 53, 14] have investigated watermark forgery under either white-box or black-box settings, where the attacker either has full access to the watermarking model or can embed watermarks into their own images. However, these approaches still rely on strong assumptions that may not hold in realistic deployment scenarios.
In contrast, the no-box setting assumes that only watermarked images are available to the attacker, without access to the model or embedding process. Yang et al. [10] proposed a heuristic method under this setting by estimating the watermark signal through averaging the residuals between watermarked and clean images, and subsequently re-embedding the estimated pattern at the pixel level. This is the scenario we focus on in this work, as it more accurately reflects practical constraints.
@@ 7 Defense Analysis
To enhance the deployed watermarking system, we suggest modifying the existing watermark system by disrupting the ability of diffusion models to model the watermark distribution effectively. Specifically, we propose a multi-message strategy as a simple yet effective countermeasure. Instead of embedding a fixed watermark message, the system randomly selects one from a predefined message pool $m_{1},m_{2},m_{3},\dots,m_{K}$ for each image. During detection, the detector verifies the presence of any valid message in the pool. This strategy introduces uncertainty into the watermark signal, increasing the entropy of possible watermark patterns and making it substantially more difficult for generative models to learn consistent features necessary for forgery. We implement this defense using different message pool sizes ($K=10,50,100$) and test on 100 images for simplicity.
As shown in the Table 4, increasing the value of $K$ leads to the FPR drops to 0% at $K=50$ and $K=100$. We further strengthen our attack by collecting more watermarked images. Specifically, we collect 5,000, 20,000, and 50,000 watermarked samples to evaluate the effect of data volume on this defense. As shown in Table 12, the FPR remained consistently at 0% even as the size of $D_{\mathrm{aux}}$ increased. Therefore, embedding multiple messages proves to be a simple yet effective countermeasure against our attack.
Table 4: Performance comparison across different $K$ values.
K=10 K=50 K=100
[…]
MS-COCO 34.73 81.63% 34.00% 34.62 69.78% 0.00% 34.86 71.56% 0.00%
CelebAHQ 36.13 83.41% 44.00% 35.89 71.00% 0.00% 35.87 72.91% 0.00%
ImageNet 34.55 79.25% 25.00% 34.35 70.09% 0.00% 34.58 71.44% 0.00%
Diffusiondb 35.14 76.28% 17.00% 35.10 70.66% 0.00% 35.40 72.28% 0.00%
Dataset 5000 20000 50000
[…]
MS-COCO 34.86 71.56% 0.00% 34.78 71.91% 0.00% 30.77 71.94% 0.00%
CelebA-HQ 35.87 72.91% 0.00% 34.15 72.97% 1.00% 27.99 72.72% 1.00%
ImageNet 34.58 71.44% 0.00% 34.57 72.56% 0.00% 30.47 72.19% 0.00%
DiffusionDB 35.40 72.28% 0.00% 34.99 72.34% 0.00% 31.15 72.06% 0.00%